In physics, the Ideal Gas Law describes the relationship among the pressure, volume, and temperature of a gas sample. This law is represented by the formula PV = nRT, where P is the pressure, V is the volume, T is the temperature, n is the amount of gas, and R is a physical constant. Which of the equations below are equivalent to the formula PV = nRT? Select all that apply.
A. P = VnRT
B. V = nRT/P
C. n = PV/RT
D. R = PVnT
E. T = nR/PV

Answer by Alan3354(69443)

In physics, --- also on most planets --- the Ideal Gas Law describes the relationship among the pressure, volume, and temperature of a gas sample. This law is represented by the formula PV = nRT, where P is the pressure, V is the volume, T is the temperature, n is the amount of gas, and R is a physical constant. Which of the equations below are equivalent to the formula PV = nRT? Select all that apply.
A. P = VnRT
Divide by V
----
B. V = nRT/P
Multiply by P
---
C. n = PV/RT
Multiply by RT
---
D. R = PVnT
E. T = nR/PV
Multiply by PV